OpenStudy (anonymous):

A 10% salt solution is to be mixed with a 20% salt solution to obtain 15 gallons of a 12% salt solution. how many gallons of each should be used?

OpenStudy (whpalmer4):

the trick with these problems is to keep track of the amount of "stuff" and the amount of liquid separately. a 10% solution has 0.1 units of salt in 1 unit of solution. if we are trying to make 15 gallons of 12% salt solution, there will be 15*0.12 = 1.8 units of salt in the 15 gallons.

OpenStudy (whpalmer4):

so we have two equations to write: let's call x the amount of 10% salt solution let's call y the amount of 20% salt solution x+y = 15 — we are making 15 gallons of 12% solution 0.1x + 0.2y = 15*0.12 — the amount of salt in the 10% solution + the amount of salt in the 20% solution = the amount of salt in 15 gallons of 12% solution

OpenStudy (whpalmer4):

it's really a coin problem in disguise :-)

OpenStudy (anonymous):

kay let me try and solve it :)

OpenStudy (anonymous):

yeah it seems almost exactly the same

OpenStudy (whpalmer4):

what do you get for your result?

OpenStudy (anonymous):

y=3? x=12?

OpenStudy (anonymous):

12+3=15

OpenStudy (whpalmer4):

excellent! you can check your answer: 12 gallons of 10% = 1.2 units of salt 3 gallons of 20% = 0.6 units of salt total 1.8 units of salt 15 gallons of 12% = 15*0.12 = 1.8 units of salt 12+3 = 15 everything checks out! just be sure to label the answer so there's no question about whether x is the 10% or the 20%

OpenStudy (whpalmer4):

heartbreaking to do the problem right and then get dinged for something stupid like switching the numbers when writing out the answer! :-)

OpenStudy (anonymous):

and i forgot my units, dang. thanks :)